OpenStudy (anonymous):

What is the exact value of the expression 2 radical 7 + radical 4 + 4 radical 7 ?

OpenStudy (lgbasallote):

dont you mean simplified? if you want exact a calculator's what you need

OpenStudy (anonymous):

ya i meant simplified

OpenStudy (anonymous):

the answer choices are: radical 28 7 radical 28 4 radical 7 + radical 14 6 radical 7 + radical 14

OpenStudy (anonymous):

@igbasallote

OpenStudy (lgbasallote):

are you sure about the question? i think you're missing something

OpenStudy (anonymous):

well the exact question i put up is the exact question, and those are the answer choices. positive

OpenStudy (lgbasallote):

\[2\sqrt 7 + \sqrt 4 + 4 \sqrt 7?\]

OpenStudy (anonymous):

if in the question the term is radical 14 instead of radical 4 then it ll be 6radical 7+radical 14

OpenStudy (anonymous):

yes @igbasallote

OpenStudy (lgbasallote):

i have no idea how to come up with one of your choices then